OpenStudy (anonymous):

find power series (maclaurin series): integral [ cos(x^3) ] dx

OpenStudy (anonymous):

Where you stuck at?

OpenStudy (anonymous):

can you take me through it from the top?

OpenStudy (anonymous):

Maclaurin series is: \[ f(x) = \sum_{n=0}^\infty f^{(n)}(0)\frac{x^n}{n!} \]

OpenStudy (anonymous):

All you really need to find is the patter for the derivative at \(0\).

OpenStudy (anonymous):

Our function in this case is: \[ f(x) = \int \cos(x^3)~dx \]Is this correct?

OpenStudy (anonymous):

I believe all we have to do is take the Maclaurin series of the integrand and integrate it, then do a bit of simplification.

OpenStudy (anonymous):

But really, where are you stuck?